# Universal Prompt Optimizer for Safe Text-to-Image Generation

### v2

@@ 2 Related Work
Prompt engineering can be categorized into three applications for foundation models: adversarial attack Xu et al. (2022), prompt tuning, and prompt optimization. By using character-level Ebrahimi et al. (2018), word-level Garg and Ramakrishnan (2020), and sentence-level Zhao et al. (2017) perturbation on prompts, attackers can launch adversarial attacks on foundation models to mislead. Prompt tuning Jia et al. (2022) is used to transform downstream tasks into pre-training tasks through constructing templates and fine-tuning models to achieve few-shot learning. Prompt optimization aims to optimize the prompt to improve the performance of prompt-based models  Hao et al. (2022); Betker et al. (2023). Prompt optimization has shown its efficiency and effectiveness in enhancing the capabilities of foundation models. In this work, we study a novel problem of prompt optimizer for T2I models to generate safe images.
@@ 3 Proposed Framework
Our prompt adaptation framework aims to produce safe prompts for T2I generation by automatically executing prompt engineering. After a user inputs a toxic prompt for the T2I generation, POSI automatically outputs the modified prompt to avoid generating inappropriate images while preserving the appropriate portion of the user’s prompt (i.e., maintaining text alignment). An illustration of the proposed framework is shown in Figure 2. Due to the absence of a publicly available toxic-clean prompts pair dataset, we first produce a set of toxic-clean prompt pairs in Section 3.1. Then we use them to conduct supervised fine-tuning (SFT) in Section 3.2 to give the model the basic ability to turn toxic prompts into clean prompts. SFT can be considered a warm-up phase, hence the effectiveness of the supervised fine-tuned model is generally moderate. To enhance the model’s performance, we further perform proximal policy optimization in Section 3.4 to maximize the target reward we design in Section 3.3, which reduces the inappropriateness of the generated images while maintaining the text alignment. Next, we give the details.
@@ 3.3 Reward Score
In order for the prompt optimizer to have the ability of rewriting toxic prompts that can generate safe and semantic-preserving images, we need to define the reward based on the modified prompt for PPO. Specifically, we measure the quality of modified prompts from two aspects: toxicity and text alignment, where toxicity measures the probability of generated images containing inappropriate content, and text alignment measures the similarity of the generated images to the text itself.
@@
To ensure that the images generated by the modified prompts still have text alignment with the original prompts, we need to quantify the relevance between the images and the original input prompts. Specifically, we compute the CLIP Radford et al. (2021) similarity scores to measure how relevant the generated images conditioned on modified prompt $x^{\prime}$ and the original input prompt $x$ are. The resulting alignment score is defined as:
@@
We use $\pi_{\phi}$ to denote the policy model to be trained during Reinforcement Learning (RL) training and $\pi_{SFT}$ to denote the supervised fine-tuned model in Section 3.2. To mitigate over-optimization Ouyang et al. (2022), we also introduce an additional KL penalty term between $\pi_{\phi}$ and $\pi_{SFT}$ with coefficient $\beta$. This is to prevent the policy model from producing meaningless prompts in pursuit of higher rewards.
@@ 3.4 Proximal Policy Optimization
With the reward score measuring both toxicity and text alignment, we propose to enhance our model by employing PPO Schulman et al. (2017) during RL trainingfor two reasons: (i) PPO been empirically proven to be data-efficient and to provide reliable performance; and (ii) We could compute the reward directly from the images produced by the T2I model to conduct PPO, without requiring knowledge of the T2I model’s internal architecture. Specifically, we initialize the parameters of $\pi_{\phi}$ by $\pi_{SFT}$. We then optimize $\pi_{\phi}$ by optimizing the following objective function in RL training over the training set $D_{PPO}=\{x\}$ as:
@@ 4.2 Baseline Methods
To test the results on SD, we choose several state-of-the-art concept removal methods. For guidance-based methods, we choose SLD Schramowski et al. (2023) under 4 settings (SLD-Weak, SLD-Medium, SLD-Strong, and SLD-Max) and SD-NP Rombach et al. (2022). For fine-tuning-based methods, we choose ESD Gandikota et al. (2023). For ESD, we fine-tune the non-cross-attention with negative guidance set as 1 (ESD-u-1). We use the prompt “hate, harassment, violence, suffering, humiliation, harm, suicide, sexual, nudity, bodily fluids, blood, obscene gestures, illegal activity, drug use, theft, vandalism, weapons, child abuse, brutality, cruelty” for SD-NP and training ESD-u-1. Note that ESD is implemented in SD v1.4, so we only present the results of ESD on SD v1.4. The implementations of all baselines are based on their official codes.
@@ 4.3 Settings
As for our framework, for the language model, we use LLaMA Touvron et al. (2023) with 7B parameters. When calculating the alignment score, we use the CLIP based on ViT-B/32. We set the T2I model $G$ in Eq. 2 as SD v1.4. We use LoRA Hu et al. (2022) for training in both the SFT and PPO phases. We set the LoRA target as q_proj and v_proj. LoRA rank is set to 8. During SFT phase, we use a batch size of 4, a learning rate of 5e-5, and a gradient accumulation step of 4. We fine-tune the model for 3 epochs (636 steps in total). During PPO phase, we use a batch size of 4, a learning rate of 1.9e-5, and a gradient accumulation step of 4. We set the $\beta$ in Eq. 4 to 0.02. We fine-tune the model for 1 epoch (52 steps in total). We generate 10 images per prompt and compute the average reward to reduce variance.
@@ 4.4 Evaluation Metric
We generate 10 images for each prompt and then we consider three evaluation metrics: (i) Inappropriate Probability (IP). We first follow the approach in Schramowski et al. (2023) to evaluate the probability of inappropriate images by combining two classifiers: the Q16 classifier Schramowski et al. (2022) and NudeNet [^2] https://github.com/notAI-tech/NudeNet. In this paper, we only consider exposed buttocks, exposed breast, exposed genitalia, exposed anus, exposed armpits, and exposed belly when using NudeNet. An image is classified as inappropriate if one or both of the classifiers output the respective label. Because our PPO phase’s reward references the output of Q16, to make a fairer comparison, we also adopt the Multi-Headed Safety Classifier (MHSC) Qu et al. (2023) as an additional classifier. We use these two methods separately to evaluate IP in the generated images; (ii) Confidence Score (CS): We then further evaluated the Confidence Score that the generated images were classified as inappropriate by the Q16 classifier; (iii) BLIP Similarity. BLIP Li et al. (2022) is an image captioning model having an image encoder and a text encoder, and BLIP similarity is calculated based on image embedding and text embedding. We use BLIP similarity between the generated images and the original prompts to evaluate the text alignment.

### v3

@@ 2 Related Work
Prompt engineering can be categorized into three applications for foundation models: adversarial attack Xu et al. (2022), prompt tuning, and prompt optimization. By using character-level Ebrahimi et al. (2018), word-level Garg and Ramakrishnan (2020), and sentence-level Zhao et al. (2017) perturbation on prompts, attackers can launch adversarial attacks on foundation models to mislead. Prompt tuning Jia et al. (2022) is used to transform downstream tasks into pre-training tasks through constructing templates and fine-tuning models to achieve few-shot learning. Prompt optimization aims to optimize the prompt to improve the performance of prompt-based models  Hao et al. (2022); Betker et al. (2023). For example, Promptist Hao et al. (2022) trains a language model to optimize user input to generate human-preferred prompts. Prompt optimization has shown its efficiency and effectiveness in enhancing the capabilities of foundation models. In this work, we study a novel problem of prompt optimizer for T2I models to generate safe images.
@@ 3 Proposed Framework
Our prompt adaptation framework aims to produce safe prompts for T2I generation by automatically optimizing prompts. Our framework is inspired by Promptist Hao et al. (2022). After a user inputs a toxic prompt for the T2I generation, POSI automatically outputs the modified prompt to avoid generating inappropriate images while preserving the appropriate portion of the user’s prompt (i.e., maintaining text alignment). An illustration of the proposed framework is shown in Figure 2. Due to the absence of a publicly available toxic-clean prompts pair dataset, we first produce a set of toxic-clean prompt pairs in Section 3.1. Then we use them to conduct supervised fine-tuning (SFT) in Section 3.2 to give the model the basic ability to turn toxic prompts into clean prompts. SFT can be considered a warm-up phase, hence the effectiveness of the supervised fine-tuned model is generally moderate. To enhance the model’s performance, we further perform proximal policy optimization in Section 3.4 to maximize the target reward we design in Section 3.3, which reduces the inappropriateness of the generated images while maintaining the text alignment. Next, we give the details.
@@ 3.3 Reward Score
In order for the prompt optimizer to have the ability of rewriting toxic prompts that can generate safe and semantic-preserving images, we need to define the reward based on the modified prompt for PPO. Specifically, the modified prompts are evaluated from two aspects: toxicity and text alignment, where toxicity measures the probability of generated images containing inappropriate content, and text alignment measures the similarity of the generated images to the text itself.
@@
To ensure that the images generated by the modified prompts still have text alignment with the original prompts, we need to quantify the relevance between the images and the original input prompts. Specifically, similar to Hao et al. (2022), we calculate the CLIP Radford et al. (2021) scores to measure how close the generated images conditioned on modified prompt $x^{\prime}$ and the original input prompt $x$ are. The alignment score is defined as:
@@
We use $\pi_{\phi}$ to denote the policy model to be trained during Reinforcement Learning (RL) training and $\pi_{SFT}$ to denote the supervised fine-tuned model in Section 3.2. To mitigate over-optimization Ouyang et al. (2022); Hao et al. (2022), we also introduce an additional KL penalty term between $\pi_{\phi}$ and $\pi_{SFT}$ with coefficient $\beta$. This is to prevent the policy model from producing meaningless prompts in pursuit of higher rewards.
@@ 3.4 Proximal Policy Optimization
With the reward score measuring both toxicity and text alignment, following Promptist Hao et al. (2022), we propose to enhance our model by employing PPO Schulman et al. (2017) during RL training for two reasons: (i) PPO has been empirically proven to be data-efficient and to provide reliable performance Schulman et al. (2017); Hao et al. (2022); and (ii) We could compute the reward directly from the images produced by the T2I model to conduct PPO, without requiring knowledge of the T2I model’s internal architecture. Specifically, we initialize the parameters of $\pi_{\phi}$ by $\pi_{SFT}$. We then optimize $\pi_{\phi}$ by optimizing the following objective function in RL training over the training set $D_{PPO}=\{x\}$ as:
@@ 4.2 Baseline Methods
To test the results on SD, we choose several state-of-the-art concept removal methods. For guidance-based methods, we choose SLD Schramowski et al. (2023) under 4 settings (SLD-Weak, SLD-Medium, SLD-Strong, and SLD-Max) and SD-NP Rombach et al. (2022). For fine-tuning-based methods, we choose ESD Gandikota et al. (2023). For ESD, we fine-tune the non-cross-attention with negative guidance set as 1 (ESD-u-1). We use the prompt “hate, harassment, violence, suffering, humiliation, harm, suicide, sexual, nudity, bodily fluids, blood, obscene gestures, illegal activity, drug use, theft, vandalism, weapons, child abuse, brutality, cruelty” Schramowski et al. (2023) for SD-NP and training ESD-u-1. Note that ESD is implemented in SD v1.4, so we only present the results of ESD on SD v1.4. The implementations of all baselines are based on their official codes.
@@ 4.3 Settings
As for our framework, for the language model, we use LLaMA Touvron et al. (2023) with 7B parameters. When calculating the alignment score, we use the CLIP based on ViT-B/32. We set the T2I model $G$ in Eq. 2 as SD v1.4. We use LoRA Hu et al. (2022) for training in both the SFT and PPO phases. We set the LoRA target as q_proj and v_proj. LoRA rank is set to 8. During SFT phase, we use a batch size of 4, a learning rate of 5e-5, and a gradient accumulation step of 4. We fine-tune the model for 3 epochs (636 steps in total). During PPO phase, we use a batch size of 4, a learning rate of 1.9e-5, and a gradient accumulation step of 4. We set the $\beta$ in Eq. 4 to 0.02. We fine-tune the model for 1 epoch (52 steps in total). Similar to Prompist Hao et al. (2022), we generate 10 images per prompt and compute the average reward to reduce variance.
@@ 4.4 Evaluation Metric
We generate 10 images for each prompt and then we consider three evaluation metrics: (i) Inappropriate Probability (IP). We first follow the approach in Schramowski et al. (2023) to evaluate the probability of inappropriate images by combining two classifiers: the Q16 classifier Schramowski et al. (2022) and NudeNet [^2] https://github.com/notAI-tech/NudeNet. In this paper, we only consider exposed buttocks, exposed breast, exposed genitalia, exposed anus, exposed armpits, and exposed belly when using NudeNet. An image is classified as inappropriate if one or both of the classifiers output the respective label as in SLD Schramowski et al. (2023). Because our PPO phase’s reward references the output of Q16, to make a fairer comparison, we also adopt the Multi-Headed Safety Classifier (MHSC) Qu et al. (2023) as an additional classifier. We use these two methods separately to evaluate IP in the generated images; (ii) Confidence Score (CS): We then further evaluated the Confidence Score that the generated images were classified as inappropriate by the Q16 classifier; (iii) BLIP Similarity. BLIP Li et al. (2022) is an image captioning model having an image encoder and a text encoder, and BLIP similarity is calculated based on image embedding and text embedding. We use BLIP similarity between the generated images and the original prompts to evaluate the text alignment.